OpenStudy (anonymous):

140 mi in 3.5 h a. 40 mi per h c. 0.03 mi per h b. 143.5 mi per h d. 35 mi per h

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

140/3.5 = 40 mil/hr

OpenStudy (anonymous):

140 mi in 3.5 h a. 40 mi per h c. 0.03 mi per h b. 143.5 mi per h d. 35 mi per h

OpenStudy (anonymous):

a

OpenStudy (anonymous):

30 mi / h = ____ ft / min a. 2640 c. 880 b. 264 d. 2112

OpenStudy (anonymous):

OMG I love u ur awesome please help me with the one up there

OpenStudy (anonymous):

The cost of printing a flyer to spread knowledge of an organization that you run is a flat rate of $9 plus $0.05 per copy. Let C represent the total cost of printing and x represent the number of copies. Write in symbolic form an equation that will relate C and x. Organize the data into a table of values. Begin with 500 copies, increase by increments of 500, and end with 2500 copies. How much would it cost to print 2500 flyers and what equation would you use? a. C = 9 + 0.05x 2500 flyers will cost $159 c. C = 9.05x 2500 flyers will cost $27,150 b. C = 9 + 0.05x 2500 flyers will cost $134 d. C = 9.05x 2500 flyers will cost $22,625